Paul's dog
Tobi weighs 95 pounds 1 pound equals 0.455 kilograms Define Toby's weight in kilograms what conversion factor should you use explain why multiplying and quantity by a conversion factor does not change the quantities value find Toby's weight in kilograms

Answer by ikleyn(52908)

.
Paul's dog
Tobi weighs 95 pounds 1 pound equals 0.455 kilograms Define Toby's weight in kilograms what conversion factor should you use
explain why multiplying and quantity by a conversion factor does not change the quantities value find Toby's weight in kilograms
~~~~~~~~~~~~~~~~~

1 pound =    0.455 kilograms.


2 pounds = 2*0.455 kilograms.


3 pounds = 3*0.455 kilograms.


. . . . . . . . 


95 pounds = 95*0.455 kilograms.


Use your calculator and calculate.


To get weight in kilograms, you need to multiply the weight in pounds by the conversion factor 0.455.